2 (absolute value bar) 3x-1 (end of absolute vale bar) +5=33
thank you!

Answer by stanbon(48551)

2|3x-1|+5 = 33
-----------------------
2|3x-1| = 28
-------
|3x-1| = 14
3x-1 = 14 or 3x-1 = -14
3x = 15 or 3x = -13
x = 5 or x = -13/3
===========================================================
Cheers,
Stan H.